Discuss "The Use of Carbon-Carbon Bond Formation in Green Chemistry" giving at least two examples. Your explanation should include relevant mechanisms and references to actual uses in life.
Answer #1

1. Formation of addition polymers.

Example: Polythene involves ethylene as the monomer and whose units are joined together to form several C-C single bonds without any formation of by-products, i.e.green chemistry pathway.

2. Diels-Alder reaction: Formation of diene-dienophile cycloaddition adducts.

Example: 1,4-butadiene reacts with ethylene to form cyclohexene as the [4+2]-cycloaddition adduct, in a green reaction pathway.

CH2=CH-CH=CH2 + CH2=CH2 ----> cyclohexene